neshia4420:

Martha earned a gross pay of $1,215.60 last week. Using the fact that Social Security tax is 6.2% of gross pay, determine the amount of Social security tax deducted from Martha’s gross pay. a. $7.54 b. $65.37 c. $75.37 d. $75.40 Please select the best answer from the choices provided A B C D

Trios:

Martha is welathy

supie:

You need to get the social security tax, I think you have to do \(1215.60 * 0.062=?\) because it says `Social Security tax is 6.2% of gross pay` So solve \(1215.60 * 0.062=?\) then you'll have your answer.

jhonyy9:

like a first step you need calcule the 6,2% of $1215,60 so for this you can use the rule of 3 simple and write in this way 1215,60 mean 100% x mean 6,2% and now cross multiplie to calcule the x value x = 6,2*1215,60/100 = ? and so in this way will get the x value what will be the correct answer hope helped understandably easy